OpenStudy (anonymous):

Four different numbers a,b,c,and d are chosen from the set {-1,-2,-3,-4,-5} .Find the largest possible value for the expression

OpenStudy (anonymous):

\[\huge a^b+c^d\]

OpenStudy (asnaseer):

hmmm... interesting problem. my first thought is that b and d must be even so that both \(a^b\) and \(c^d\) are positive

OpenStudy (asnaseer):

so then problem reduces to finding maximum value of this expression:\[\frac{1}{a^2}+\frac{1}{b^4}\]with a and b drawn from the set {-1, -3, -5}

OpenStudy (shubhamsrg):

naturally we should go for -1 and -3.

OpenStudy (shubhamsrg):

-1 = b and -3 =a

OpenStudy (asnaseer):

exactly

OpenStudy (asnaseer):

I was /hoping/ that @Jonask would fill in the blanks :)

OpenStudy (anonymous):

lol yes you made a good start a=-3 and b=-1 since \[\left| a \right| \ge \left| b \right|\]

OpenStudy (anonymous):

5 would yield a smaller result so not part of the solution

OpenStudy (asnaseer):

yup :)